$SPAD/src/input f01ref.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)cl all
showArrayValues true
showScalarValues true
wheret:="s"
m:=5
n:=3
ncolq:=2
lda:=5
theta:Matrix Complex SF:=
 [[1 + 0.5*%i ,1.0954-0.3333*%i ,1.2649-1.1565*%i ]]
a:Matrix Complex SF :=
 [[ 1             , 1      +1     *%i , 1     + %i ],_
  [-0.2  - 0.4*%i ,-2                 ,-1     - %i ],_
  [-0.32 - 0.16*%i,-0.3505 +0.2629*%i ,-3          ],_
  [-0.4  + 0.2*%i ,         0.5477*%i , 0.0        ],_
  [-0.12 + 0.24*%i, 0.1972 +0.2629*%i , 0.6325     ]]
result:=f01ref(wheret,m,n,ncolq,lda,theta,a,-1)

———-

References^{[1] nothing}

4